OpenStudy (anonymous):

How do you simplify this and find the domain? e^logx/(x + 8)

OpenStudy (anonymous):

\[\frac{e ^{\log(x)}}{x+8}?\]

OpenStudy (anonymous):

no the x + 8 where you have the x beside log

OpenStudy (anonymous):

use the equation on the bottom rewrite

OpenStudy (sasogeek):

\[\huge e^{{log\frac{x}{x+8}}} \]

OpenStudy (sasogeek):

is that the question?

OpenStudy (anonymous):

yes

OpenStudy (campbell_st):

well e^log(x) = x so in your question it simplifies to x/(x+8) as far ad the domain is concerns x cannot equal 8... but can be any other real values... so domain, all real x except x=8

OpenStudy (sasogeek):

cambell, \[\huge e^{log(x)} \neq x \ but \ rather, e^{ln(x)}=e^{\log _{e}x=x}\] \]

OpenStudy (campbell_st):

well use change of base to change it to base e logs the domain is still unchanged.

OpenStudy (campbell_st):

so ln (x/x+8) = ln(x/x+8)/ln10 thats change of base...

OpenStudy (campbell_st):

then the problem is e^ln(etc)...

OpenStudy (campbell_st):

simplfies to e^ln(x/10(x+8) = x/10(x+8) domain is unchanged.

OpenStudy (sasogeek):

\[\huge \text {taking only the exponent...} \]\[\huge log(\frac{x}{x+8})=\frac{ln(\frac{x}{x+8})}{ln(10)}=\] ....

OpenStudy (sasogeek):

how is it \[\huge \frac{x}{10(x+8)} ? \]

OpenStudy (campbell_st):

well log (x/(x+8) -10) then the domain is all real x except x cannot equal 2...

OpenStudy (sasogeek):

well the questions says simplify... not solve... :P

OpenStudy (campbell_st):

it says state the domain

OpenStudy (anonymous):

logx/x + 8 = Inx/(x + 8)In10?

OpenStudy (sasogeek):

change of base

OpenStudy (anonymous):

log(4)/log(e) != 4/e

OpenStudy (anonymous):

Can't cancel out functional operators as with dviison.

OpenStudy (sasogeek):

well i'm lost here so i'll delete my solutions, if i find something i can settle with as a solution and i'm sure of it, i'll let u guys know, at the moment, i'm still confused about all the current answers but if that is the right solution, awesome! i'm the confused one here :)

OpenStudy (anonymous):

The simplification I have is (x^(1/ln(10)))/((x+8)^(1/ln(10)))

OpenStudy (anonymous):

What the hell is it with transcendental numbers.

OpenStudy (anonymous):

Alright, because the exponent is a transcendental number, I'm going to say that x must be x>0 for the top. For the bottom, (x+8) must be (x+8)>0, or x>-8.

OpenStudy (sasogeek):

\[\huge e^{ln\frac {(\frac{x}{x+8})}{ln10}}=\frac{(ln10)x}{x+8} \]

OpenStudy (anonymous):

Continued discussion : http://openstudy.com/study#/updates/4f38ae13e4b0fc0c1a0dda3b